OpenStudy (anonymous):

What is the answer to this problem? (Simplify) 11x − (2x2 + 8x).

OpenStudy (anonymous):

2x^2

OpenStudy (anonymous):

11x-2x^2-8x=0 3x=2x^2 x^2=3/2 x=sqrt3/sqrt2 calculate the rest

OpenStudy (anonymous):

I'm not exactly sure that is correct sir.

OpenStudy (aravindg):

hii

OpenStudy (aravindg):

hope u r satisfied!!giv medal lol